# Conformal Online Learning of Deep Koopman Linear Embeddings

### v1

@@ B.1.1 Simulated data
This system, studied in depth in Section 4.1 and also considered in Section 4.2, is a simple 2D ODE whose dynamic converges to a known attracting manifold [Brunton2016Control, Bruton2022ModernKoopman]. Trajectories are simulated via odeint from SciPy with 100 time steps, equally spaced with step size $\Delta t=0.01$, using initial conditions sampled uniformly from the domain $2,2]^{2}$. The continuous-time eigenvalues reported in Section 4.1 are obtained from the estimated discrete-time ones via the transformation $\lambda=\log(\lambda_{\text{disc}})/\Delta t$. The ground-truth Koopman eigenfunctions for this system are given by $\varphi_{1}=x_{2}-\frac{\lambda_{1}^{*}}{\lambda_{1}^{*}-2\lambda_{2}^{*}}\cdot x_{1}^{2}$, $\varphi_{2}=x_{1}$, and $\varphi_{3}=x_{1}^{2}$.
@@
where the parameters $\delta$, $\beta$, and $\mu$ represent the damping coefficient, the linear stiffness, and the nonlinear stiffness respectively. Here, we choose the parameters $\delta=0.5$, $\beta=-1$, and $\mu=1$, following the setting used in [Williams2015EDMD, Otto2019Autoencoder]. This nonlinear system exhibits two stable spirals at $u=\pm 1,\dot{u}=0$ and a saddle at the origin. Trajectories are simulated using odeint, each consisting of 100 time steps with a fixed interval $\Delta t=0.025$. Initial conditions are sampled uniformly from the domain $2,2]^{2}$.
@@
where $\mu>0$ controls the nonlinearity and damping strength. We follow the setup in sinha2019rRecursiveEDMD and set $\mu=0.2$. Trajectories are simulated using odeint, with 100 time steps and a fixed step size of $\Delta t=0.01$. Initial conditions are sampled uniformly from the square domain $4,4]^{2}$.
@@
where we have chosen $\sigma=10$, $\rho=28$, and $\beta=8/3$, a commonly studied parameter regime known to induce chaotic behavior which was also considered in [kostic2022rkhs]. We simulate the trajectories using RK45, with a time step $\Delta t=0.01$ over 500 steps. Initial conditions are sampled uniformly from the cube $10,10]^{3}$ centered at the origin.
@@ B.1.2 Real data
We test on ETTh1 which is part of the Electricity Transformer Dataset (ETDataset) [^0] https://github.com/zhouhaoyi/ETDataset introduced by zhou2021informer. ETTh1 contains data recorded at hourly intervals for roughly two years from an electricity transformer station. The dataset consists of a single trajectory of 6 features: HUFL (High Use Frequency Load), HULL (High Use Low Load), MUFL (Medium Use Frequency Load), MULL (Medium Use Low Load), LUFL (Low Use Frequency Load), and LULL (Low Use Low Load). In our setting, we retain 200 time steps and aim to learn the transformer’s load profile as a dynamical system.
@@
For all datasets, data prior to time $t_{0}$ are used to initialize the model parameters, including those of the Conformal PI control for COLoKe [Angelopoulos2023ConformalPID]. In the case of synthetic datasets, we set $t_{0}=T/5$, resulting in $t_{0}=20$ for the single attractor, Duffing oscillator, and Van der Pol oscillator, and $t_{0}=100$ for the Lorenz system. For all real datasets, we use $t_{0}=100$.
@@ B.2 Baseline Models
Online Dynamic Mode Decomposition [Zhang2018OnlineDMD] incrementally estimates a linear model from sequential data, enabling efficient updates without storing the entire dataset. Given a trajectory $\{x_{1},x_{2},\dots x_{t_{0}}\}$ available to time $t_{0}$ for model initialization. Let $X_{t_{0}}=\left[x_{1},x_{2},\dots,x_{t_{0}-1}\right]$ and $Y_{t_{0}}=\left[x_{2},x_{3},\dots,x_{t_{0}}\right]$. The matrix $K_{t_{0}}$ is computed by
@@
This online formulation allows for recursive updates of $K_{t}$ and yields an efficient and memory-light algorithm for learning linear dynamics in real time. The readers are referred to [Zhang2018OnlineDMD] for a comprehensive explication of ODMD.
@@
We first present Recursive EDMD [sinha2019rRecursiveEDMD] and point out its limitations for pure online setting, in order to introduce the variant OEDMD. Choose a fixed dictionary $\Phi:\mathcal{X}\to\mathbb{R}^{r}$
@@
In the original work [sinha2019rRecursiveEDMD], the authors used Radial Basis Functions (RBF) as the fixed dictionary. However, to estimate informative centers for RBF, one needs to have access to the full trajectory up to time $T$. When estimating the centers with trajectory up to time $t_{0}$, the model gives poor results on all datasets for both metrics. Therefore, we choose a polynomial dictionary of degree 2, which provides a lifted representation dimension comparable to other baseline models. Increasing the degree beyond 2 offers no significant performance gain. On the contrary, it degrades the performance on the real dataset and results in substantially higher computational costs.
@@
The detailed presentation of COLoKe can be found in Section 3. The neural network $\tilde{\Phi}$ is fully connected with architecture $\{d,32,16,8,m-d\}$ for synthetic datasets and $\{d,64,32,16,m-d\}$ for the real dataset. The dimension of lifted representation $m$ is chosen to be $d+\left\lceil d/2\right\rceil$ for all experiments. Model parameters and initial conformity threshold are initialized with $\{x_{0},\cdots,x_{t_{0}}\}$ as already discussed. The initial threshold $q_{t_{0}+1}$ is set to be $1-\alpha$ quantile of the set of scores $\{s_{w+1},\dots,s_{t_{0}}\}$ computed with the initialized model. The model parameters $(\theta_{t},K_{t})$ are updated online according to Algorithm 1. For all synthetic datasets, the hyperparameters for Conformal PI procedure are $\alpha=0.5,\gamma=0.1,C_{sat}=5$, and we set $C_{sat}=10$ for the real dataset. Note that the coverage guarantees [Angelopoulos2023ConformalPID] hold for any value of $\gamma>0$. We set $\gamma=0.1$ in our experiments as in [Angelopoulos2023ConformalPID], although one may implement adaptive strategies as in [bhatnagar2023improved]. The choice of $\alpha$ should reflect the trade-off between computational efficiency and predictive accuracy, depending on the requirements of the application. When $\alpha\approx 1$, accuracy is prioritized over speed, whereas for $\alpha\ll 1$, computational speed is more important. In our experiments, we selected $\alpha=0.5$ as a balanced compromise between these two objectives.
@@
We implement the model in [liang2022online]. The original work tackles the problem of Model Predictive Control (MPC). We aim only to perform online learning of dynamics. For synthetic datasets, the encoder architecture is $\{d,32,16,8,m\}$ and the decoder architecture is $\{m,8,16,32,d\}$. For the real dataset, the encoder architecture is $\{d,64,32,16,m\}$ and the decoder architecture is $\{m,64,32,16,d\}$. The dimension of lifted representation m is chosen to be $d+\left\lceil d/2\right\rceil$. Thus, the model architecture of OnlineAE aligns with COLoKe and OLoKe. The loss function at time $t$ is defined as

### v2

@@ B.1.1 Simulated data
This system, studied in depth in Section 4.1 and also considered in Section 4.2, is a simple 2D ODE whose dynamic converges to a known attracting manifold [Brunton et al., 2016, 2022]. Trajectories are simulated via odeint from SciPy with 100 time steps, equally spaced with step size $\Delta t=0.01$, using initial conditions sampled uniformly from the domain $2,2]^{2}$. The continuous-time eigenvalues reported in Section 4.1 are obtained from the estimated discrete-time ones via the transformation $\lambda=\log(\lambda_{\text{disc}})/\Delta t$. The ground-truth Koopman eigenfunctions for this system are given by $\varphi_{1}=x_{2}-\frac{\lambda_{1}^{*}}{\lambda_{1}^{*}-2\lambda_{2}^{*}}\cdot x_{1}^{2}$, $\varphi_{2}=x_{1}$, and $\varphi_{3}=x_{1}^{2}$.
@@
where the parameters $\delta$, $\beta$, and $\mu$ represent the damping coefficient, the linear stiffness, and the nonlinear stiffness respectively. Here, we choose the parameters $\delta=0.5$, $\beta=-1$, and $\mu=1$, following the setting used in [Williams et al., 2015, Otto and Rowley, 2019]. This nonlinear system exhibits two stable spirals at $u=\pm 1,\dot{u}=0$ and a saddle at the origin. Trajectories are simulated using odeint, each consisting of 100 time steps with a fixed interval $\Delta t=0.025$. Initial conditions are sampled uniformly from the domain $2,2]^{2}$.
@@
where $\mu>0$ controls the nonlinearity and damping strength. We follow the setup in Sinha et al. [2019] and set $\mu=0.2$. Trajectories are simulated using odeint, with 100 time steps and a fixed step size of $\Delta t=0.01$. Initial conditions are sampled uniformly from the square domain $4,4]^{2}$.
@@
where we have chosen $\sigma=10$, $\rho=28$, and $\beta=8/3$, a commonly studied parameter regime known to induce chaotic behavior which was also considered in [Kostic et al., 2022]. We simulate the trajectories using RK45, with a time step $\Delta t=0.01$ over 500 steps. Initial conditions are sampled uniformly from the cube $10,10]^{3}$ centered at the origin.
@@ B.1.2 Real data
We test on ETTh1 which is part of the Electricity Transformer Dataset (ETDataset) [^0] https://github.com/zhouhaoyi/ETDataset introduced by Zhou et al. [2021]. ETTh1 contains data recorded at hourly intervals for roughly two years from an electricity transformer station. The dataset consists of a single trajectory of 6 features: HUFL (High Use Frequency Load), HULL (High Use Low Load), MUFL (Medium Use Frequency Load), MULL (Medium Use Low Load), LUFL (Low Use Frequency Load), and LULL (Low Use Low Load). In our setting, we retain 200 time steps and aim to learn the transformer’s load profile as a dynamical system.
@@
For all datasets, data prior to time $t_{0}$ are used to initialize the model parameters, including those of the Conformal PI control for COLoKe [Angelopoulos et al., 2023]. In the case of synthetic datasets, we set $t_{0}=T/5$, resulting in $t_{0}=20$ for the single attractor, Duffing oscillator, and Van der Pol oscillator, and $t_{0}=100$ for the Lorenz system. For all real datasets, we use $t_{0}=100$.
@@ B.2 Baseline Models
Online Dynamic Mode Decomposition [Zhang et al., 2019] incrementally estimates a linear model from sequential data, enabling efficient updates without storing the entire dataset. Given a trajectory $\{x_{1},x_{2},\dots x_{t_{0}}\}$ available to time $t_{0}$ for model initialization. Let $X_{t_{0}}=\left[x_{1},x_{2},\dots,x_{t_{0}-1}\right]$ and $Y_{t_{0}}=\left[x_{2},x_{3},\dots,x_{t_{0}}\right]$. The matrix $K_{t_{0}}$ is computed by
@@
This online formulation allows for recursive updates of $K_{t}$ and yields an efficient and memory-light algorithm for learning linear dynamics in real time. The readers are referred to [Zhang et al., 2019] for a comprehensive explication of ODMD.
@@
We first present Recursive EDMD [Sinha et al., 2019] and point out its limitations for pure online setting, in order to introduce the variant OEDMD. Choose a fixed dictionary $\Phi:\mathcal{X}\to\mathbb{R}^{r}$
@@
In the original work [Sinha et al., 2019], the authors used Radial Basis Functions (RBF) as the fixed dictionary. However, to estimate informative centers for RBF, one needs to have access to the full trajectory up to time $T$. When estimating the centers with trajectory up to time $t_{0}$, the model gives poor results on all datasets for both metrics. Therefore, we choose a polynomial dictionary of degree 2, which provides a lifted representation dimension comparable to other baseline models. Increasing the degree beyond 2 offers no significant performance gain. On the contrary, it degrades the performance on the real dataset and results in substantially higher computational costs.
@@
The detailed presentation of COLoKe can be found in Section 3. The neural network $\tilde{\Phi}$ is fully connected with architecture $\{d,32,16,8,m-d\}$ for synthetic datasets and $\{d,64,32,16,m-d\}$ for the real dataset. The dimension of lifted representation $m$ is chosen to be $d+\left\lceil d/2\right\rceil$ for all experiments. Model parameters and initial conformity threshold are initialized with $\{x_{0},\cdots,x_{t_{0}}\}$ as already discussed. The initial threshold $q_{t_{0}+1}$ is set to be $1-\alpha$ quantile of the set of scores $\{s_{w+1},\dots,s_{t_{0}}\}$ computed with the initialized model. The model parameters $(\theta_{t},K_{t})$ are updated online according to Algorithm 1. For all synthetic datasets, the hyperparameters for Conformal PI procedure are $\alpha=0.5,\gamma=0.1,C_{sat}=5$, and we set $C_{sat}=10$ for the real dataset. Note that the coverage guarantees [Angelopoulos et al., 2023] hold for any value of $\gamma>0$. We set $\gamma=0.1$ in our experiments as in [Angelopoulos et al., 2023], although one may implement adaptive strategies as in [Bhatnagar et al., 2023]. The choice of $\alpha$ should reflect the trade-off between computational efficiency and predictive accuracy, depending on the requirements of the application. When $\alpha\approx 1$, accuracy is prioritized over speed, whereas for $\alpha\ll 1$, computational speed is more important. In our experiments, we selected $\alpha=0.5$ as a balanced compromise between these two objectives.
@@
We implement the model in [Liang et al., 2022]. The original work tackles the problem of Model Predictive Control (MPC). We aim only to perform online learning of dynamics. For synthetic datasets, the encoder architecture is $\{d,32,16,8,m\}$ and the decoder architecture is $\{m,8,16,32,d\}$. For the real dataset, the encoder architecture is $\{d,64,32,16,m\}$ and the decoder architecture is $\{m,64,32,16,d\}$. The dimension of lifted representation m is chosen to be $d+\left\lceil d/2\right\rceil$. Thus, the model architecture of OnlineAE aligns with COLoKe and OLoKe. The loss function at time $t$ is defined as